OpenStudy (anonymous):

log2x−log4x=−2 how then? :/

OpenStudy (anonymous):

\[\log _{2}x-\log _{4}x\]

OpenStudy (anonymous):

=-2

OpenStudy (anonymous):

\[\log _{2}(x \div 4) = \log _{2} (4x)\] and hence i got x=16x lol

OpenStudy (anonymous):

but that is wrong for sure x is not 0

OpenStudy (anonymous):

the answer is 1 over 16 btw

OpenStudy (dumbcow):

i think you have to use change of base formula \[\log_{b} x = \frac{\ln x}{\ln b} \]

OpenStudy (anonymous):

Hmm i nvr did algorithm for a long time, not sure if im right. But here it is, Do a change of base. \[\log_{2} x - \log_{4} x= -2\] \[\frac{\log_{4}x}{\log_{4}2}-\log_{4}x = -2\] \[\frac{\log_{4}x}{2} - \log_{4}x= -2\] \[\log_{4} x- 2 \log_{4}x = -4\] \[\log_{4} x = 4\] \[4= x ^{4}\] \[x=4^{\frac{1}{4}}\] x= 1.414

OpenStudy (anonymous):

well i didnt get the answer, hmm think i went wrong somewhr.

OpenStudy (anonymous):

\[\log _{4}2 \] is not 2 lol i saw ur mistake

OpenStudy (anonymous):

Oh my god. told you im rusty. i got the algorithm totally the wrong way. damn it

OpenStudy (dumbcow):

\[\rightarrow \frac{\ln x}{\ln 2} - \frac{\ln x}{\ln 4} = -2\] ln(4) = ln(2^2) = 2*ln(2) \[\rightarrow \frac{2\ln x - \ln x}{2 \ln 2} = -2\] \[\ln x = -4*\ln 2\] \[x = e^{-4 \ln 2} = 2^{-4} = \frac{1}{16}\]

OpenStudy (anonymous):

that is complicating but i get it thanks :)

OpenStudy (dumbcow):

your welcome, tiaph was correct as well...you can change it to any base you like...i just like using natural logs :)